Institut for Matematik og Datalogi Odense Universitet February 2, 1995

JFB

DM15 – F95 – Week 16

Lecture, February 6

We will cover sections 8.1 and 8.2 in Sudkamp.

Lecture, February 13

We will finish chapter 8 and begin chapter 9 in Sudkamp.

Problems to be discussed in week 6

These were on the the last weekly note for 1994.

Problems to be discussed in week 7

• Do problem 5 on page 207 of Sudkamp.

• Do problem 7 on page 207 of Sudkamp.

• Do problems 12a, 12c, and 12e on page 208 of Sudkamp.

• Do problem 17 on page 209 of Sudkamp. As the problem is stated, it is impossible to prove that L(M) = L(G), because this is false. Thus, you should first figure out what’s wrong with the PDA and how to correct it. Hint: How should it start its computation?

Note

There will only be one discussion section each week this semester. Peter Høyer will be the instructor, and it will be on Thursdays at 14:15, in U8.